A study was conducted to determine the proportion of people who dream in black and white instead of color. Among 288 people over the age of​ 55, 80 dream in black and​ white, and among 296 people under the age of​ 25, 13 dream in black and white. Use a 0.01significance level to test the claim that the proportion of people over 55 who dream in black and white is greater than the proportion for those under 25. Complete parts​ (a) through​ (c) below.

- Identify the p value?

- Test the claim by constructing an appropriate confidence level?

- What is the conclusion base on the hypothesis test?

- What is the conclusion base on the confidence level?
Answer #1

Answer: 288 Total e) 296 (P2) Bow Color Age > 55 80 208 Age 225 13 283 Hypothesis Hoip, =P2 H, P, P2 P, = 80 -0.2777 P2= 288Given Q-0.01 So confidence level=99% for this (one talled) Z = 2.326 I zxo бе. -Р, confidence interval Sample statistics (P,